# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v1

@@ VI Limitations
Our solution, FairLay-ML, currently supports tabular data in the form of a CSV and binary classification algorithms from sci-kit learn. Decision Trees, Support Vector Machines, Random Forest, and Logistic Regression models are tested. We plan to extend the tool to support datasets beyond tabular ones as well as training algorithms provided by TensorFlow and PyTorch.
VII Related Work
@@
Seldonian [7] presented a mitigation technique that allows users to directly specify (arbitrary) undesirable behaviors. The key innovation of Seldonian is to enforce safety requirements during training in the offline contextual bandits setting with probabilistic guarantees. FairLay-ML, on the other hand, focuses on helping users detect and understand discrimination. IBM AIF360 [8] and Fairlearn [9] implement various measures of fairness and algorithms to mitigate unfairness in three stages pre-processing, in-processing, and post-processing. However, they focus on group fairness and evaluate unfairness over a pre-fixed development set. Google What-if toolkit [11] is the closest to FairLay-ML, and we discussed it in detail in the introduction section. eXplainable AI (e.g., LIME [25], SHAP [32], etc.) are also relevant, and FairLay-ML significantly adapted XAI and visualization tools to enable users to debug their data-driven software for fairness. However, those tools try to explain how ML models globally (for all data points) and locally (for a specific data point) make decisions. FairLay-ML focuses specifically on fairness and allows users to identify significant test cases and compare multiple data points simultaneously.
VIII Conclusion
We presented FairLay-ML, an intuitive debugging tool that aids data-driven users to analyze and understand fairness implications of their solutions. FairLay-ML integrates visualization, testing, and human annotations to help developers navigate the complexity of data-driven solutions. One exciting future direction is to support mitigating unfairness. One idea that is partially implemented in the current tool is to mask features which are highly correlating with protected attributes. We plan to enable users to perform targeted data masking. For example, in the relationship feature in the Census data set, there are 5 categories: husband, wife, not married, married with children, and spouse in military; it may make sense to only mask husband and wife, as proxies for gender, without masking the other categories.

### v2

@@ V Limitations
Our solution, FairLay-ML, currently supports tabular data in the form of a CSV and binary classification algorithms from scikit-learn such as Decision Trees, SVMs, Random Forest, and Logistic Regression.
VI Related Work
@@
Seldonian [16] presented a mitigation technique that allows users to directly specify (arbitrary) undesirable behaviors. The key innovation of Seldonian is to enforce safety requirements during training in the offline contextual bandits setting with probabilistic guarantees. FairLay-ML, on the other hand, focuses on helping users detect and understand discrimination. While eXplainable AI (e.g., LIME [14], SHAP [17], etc.) explain how ML models globally (for all data points) and locally (for a specific data point) make decisions, FairLay-ML focuses specifically on fairness and allows users to identify significant test cases and compare data points.
VII Future Work
We presented FairLay-ML, an intuitive debugging tool to understand the implications of the fairness of sociotechnological software. One exciting future direction is to support mitigating unfairness. One idea that is partially implemented in the current tool is to mask features which are highly correlating with protected attributes.
Acknowledgement. Carreon was supported by NSF grant CNS-2137791 through CAHSI REU program. This research has been supported by NSF under Grant No. CNS-2230060 and CNS-2230061.